# Consensus formation in quality-sensitive interdependent agent systems

### v2

@@ III.2 Asymmetric discovery scenario
III.3 Finite Size Effects
@@
In the lower-left corner of Fig. 7B, we observe a null variance, followed by a sudden increase at higher values of $\lambda$. This corresponds to the region where the system ends up in a fully uncommitted state, occurring when $\pi_{1,2}=0$ and $\lambda<r_{2}$ in an infinite system. Although analytically the change in behavior is predicted to occur precisely at $\lambda=r_{2}$, we note that due to finite-size effects, this behavior extends slightly beyond this threshold. At $\lambda\gtrsim r_{2}$, fluctuations suddenly increase as most of the simulation results converge around the predicted stable solution, $f_{0}=r_{2}/\lambda$, while, due to finite-size effects, some results still show the system stabilizing at the fully uncommitted state. As the system size increases, the transition becomes sharper around the threshold $\lambda=r_{2}$.
@@
Lastly, let’s consider the scenario where the site qualities are equal (and their discovery probabilities are also equal) in the fully interdependent limit ($\pi_{\alpha}=0$). As we’ve discussed in the analytical solution of the model, detailed in Sec. II.1.2, the system may undergo a phase transition between a fully uncommitted absorbing state and an active state with finite fractions of bees advertising the available sites at $\lambda=r$. We further investigate this non-equilibrium phase transition in the following section. Here, we briefly emphasize the influence of finite-size effects under this limit. While the system may temporarily fluctuate around the stationary state predicted by the deterministic equations, finite-size effects ultimately break this tendency, and the system converges towards a consensus for only one of the options, while the other option disappears. In Figs. 8, we illustrate this phenomenon, by representing the temporal evolution of $f_{0}$, $f_{1}$ and $f_{2}$ for a binary decision problem starting from two different initial conditions. Continuous lines represent the results obtained in stochastic simulations of a finite system consisting of $N=1000$ bees, while dashed lines represent the integration of deterministic equations Eqs. (4), representing the behavior of an infinite system. We observe that fluctuations introduced by finite-size effects ultimately break the symmetry between the two available options. This can occur even when one option initially appears to be favored.
@@
In real-world applications, whether in ecological systems or synthetic ones like robot swarms, these fluctuations are always expected to be relevant because most of these systems are finite. An important conclusion can be drawn from these results. In ecological systems, the ability to regulate the dissemination of information is crucial for efficient decision-making. For instance, when options have comparable qualities, it may be beneficial for the group to converge quickly towards a consensus to avoid prolonged indecision. In such cases, certain members within the group may play a pivotal role by transmitting stop exploration signals to others [20]. These signals would effectively reduce the parameter $\pi$ to near zero, signaling to other group members to cease exploration and focus on promoting consensus for one of the options. This adaptive behavior allows the group to navigate decision-making scenarios with similar options more effectively, potentially avoiding resource wastage. However, deterministic modeling approaches, may struggle to capture the dynamics of decision-making processes in real finite systems in similar circumstances.
@@
III.4 Absorbing phase transition and finite size effects in the fully interdependent limit
@@
The self-discovery probabilities $\pi$ play the role of an external field, as with non-null $\pi$ values, an inactive site can spontaneously become active. Similar to the contact process, the introduction of this external field disrupts the absorbing state and consequently the transition itself, leading the system away from criticality. However, for sufficiently small values of $\pi$, the system still obeys certain scaling laws, as discussed in [43]. We have also confirmed that, for instance, the order parameter in stationary conditions is compatible with the scaling law $\rho\sim\pi^{\beta/\sigma}\sim\pi^{1/2}$, with mean-field exponent values $\beta=1$ and $\sigma=2$.
Refer to captionRefer to caption
@@
Non-equilibrium phase transitions are commonly observed in biological systems [46]. These transitions, which occur far from thermodynamic equilibrium, play a crucial role in shaping the collective behavior of living organisms. Here, we find another example which might be relevant for collective decision making processes.
III.5 Weber’s Law of perception
Psychophysics explores how organisms perceive external stimuli, initially concentrating on human perception before broadening its scope to encompass other organisms across varying levels of biological complexity. There’s a proposition that a swarm engaged in decision-making tasks can be seen as a superorganism [30], exhibiting cognitive traits akin to those found in vertebrate brains based on neurons. Building on recent research into honeybee-inspired decision-making models [36], we aim to ascertain whether the LES model also reflects such characteristics.
We particularly investigate Weber’s Law [47], which posits that the smallest perceptible change between two stimuli (known as the just noticeable difference) is proportional to the intensity of the base stimulus. In our context, the stimuli correspond to the qualities of the two options under consideration in a binary decision problem, and the base stimulus intensity can be defined as the mean quality of the options, denoted as $\bar{q}$. According to Weber’s Law, we anticipate observing a linear correlation between the difference in quality and the base stimulus strength ($\bar{q}$), expressed as $q_{2}-q_{1}=w\bar{q}$, where $w$ represents the Weber fraction.
Following the analysis in [36], one can also relate the magnitude of finite-size fluctuations within the decision process with the random fluctuations typically observed in a discrimination process. These fluctuations can influence an organism’s ability to accurately discriminate between two stimuli [48, 49]. It is widely recognized that collective decisions tend to improve when made by larger groups [50, 51]. As discussed in Sec. III.3, fluctuations diminish as the system size increases. Consequently, we anticipate that the system’s ability to discriminate between two similar stimuli will improve with the system size, i.e. the just noticeable difference will decrease with $N$. In agreement with [36], we observe this effect in our model, as illustrated in Fig. 11C, which also exhibits a comparable exponential trend.
In the study by Reina et al. [36], individual commitment transitions (i.e., discovery) are linked to site qualities, unlike in our model. Here, the quality-sensitivity influences system dynamics solely in the uncommitment and recruitment transitions. However, this difference does not appear to impact the model’s adherence to Weber’s Law. Ihe independent discovery probabilities solely affect the specific value of quality differences that can be discerned. Figures 11 A and B also demonstrate that, for a given stimulus strength, the hability to discriminate decreases with $\lambda$. In other words, stronger social interaction enables the system to distinguish between more similar options, but this comes at the expense of a slower decision-making process, as previously noted in Figs. 4 and 6.
Contrary to our findings, Reina et al. in [36] observe that the noticeable quality difference increases (while the decision time decreases) with the signaling ratio, which in their setup reflects the strength of social interactions relative to individual transitions. This disparity underscores the main distinction between the two models: the presence (or absence, in our case) of negative feedback in social interactions, specifically cross-inhibition. Introducing this mechanism in the decision process leads to quicker decisions, especially in situations involving similarly or equally valued options [24], albeit at the expense of reduced accuracy in the final outcome. These results underscore not only the significance of social interactions but also their nature, as the presence or absence of cross-inhibition yields markedly different outcomes in scenarios initially featuring very similar problems.
@@ III.6 Multiple sites
Up to now, our attention has been directed towards a binary decision problem, where selecting the best option is imperative. Many collective decision making models have been concerned in generalizing their dynamics to multiple option scenarios [37, 24, 39, 40]. Particularly, previous research on this model, utilizing different parameter setups, examined a specific scenario involving five options. It’s worth noting that increasing the number of available options doesn’t qualitatively alter the model’s behavior. In a general $k$-site scenario, we still observe the same interplay between independent discovery and social feedback. At high $\lambda$ values, the system effectively identifies the best site, while at low $\lambda$ values, it settles into a multi-opinion state where each site is occupied in proportion to its quality and discovery probability. Therefore, in a broader best-of-N scenario, the specific dance frequency values and consensus transitions will rely on the $\pi_{\alpha}$ and $q_{\alpha}$ parameters defining each unique scenario. However, the overall behavioral trend of the system remains consistent across different setups.
@@
As the number of available options with the same discovery probability increases, the system is exposed to a greater amount of external information. Consequently, slightly higher values of $\lambda$ are required to achieve the consensus crossover, as evidenced by the rightward shift of the consensus curves in Fig. 12A around $Q=0$. Contrarily, when the overall level of external information ($\sum\pi_{\alpha}$) remains constant but the individual discovery probabilities $\pi_{\alpha}$ decrease as the number of options increases, as shown in Fig. 12B - the system achieves consensus more readily as the number of options increases, a result that may appear counter-intuitive compared to other studies on the best-of-N scenario with different models [24]. However, this outcome is a direct consequence of the decision mechanisms embedded in our model: increasing $\lambda$ strengthens the positive feedback on populations associated with the best quality, leading to the rapid elimination of options with lower qualities from consideration. Consequently, the decision process becomes nearly binary between the top two options. Decreasing the overall values of the discovery probabilities simplifies the discussion process further, facilitating consensus. Supplementary Fig. LABEL:suppfig:all_fs_varNsites illustrates all dance frequencies $f_{\alpha}$ in the reported scenarios, demonstrating the consistently decreasing trends of the remaining populations in favor of the highest-quality option as $\lambda$ increases.
Other scenarios, such as a generally asymmetric discovery scenario (where $\pi_{1}>\pi_{2}>...>\pi_{k}$) or equivalent inferior options ($q_{1}=...=q_{k-1}$), exhibit similar trends to the discussed symmetric discovery scenario. Naturally, the specific values of $f_{k}$ or the consensus may vary, as it becomes more challenging for the swarm to achieve consensus under these conditions. However, the interplay between interdependence, site qualities, and discovery probabilities remains unchanged. In Supplementary Fig. LABEL:suppfig:moreSitesDif we present the results obtained in these scenarios to emphasize the qualitative similarities discussed earlier.
@@ IV Conclusions
We have performed an analytical study of an agent-based model inspired by the collective behavior of honeybees [21], following the mathematical framework outlined in [31]. This model incorporates key features such as the balance between independent exploration and social interdependence, as well as the sensitivity of agents to option quality. Using analytical reductions of this agent-based model, we extensively analyzed its behavior in the space of relevant parameters. Our goal was to investigate whether the system achieves consensus for the best available option when balancing information acquisition and social interactions. Furthermore, we supplemented this analysis with simulations to report the time required for the system to converge to its stationary state.
@@
We have explored two scenarios: one where options are equally likely to be discovered, and another where the worst quality option (among two) is more likely to be discovered. In both cases, when the system prioritizes interdependence, the established consensus is stronger. This highlights the role of social interactions in achieving a higher accuracy in the final decision, consistent with similar honeybee-inspired models [23, 24, 25]. Interdependence plays a critical role, especially when the system deals with high values of the self-discovery probabilities or when the superior option faces a disadvantage in terms of discovery likelihood. In such scenarios, interdependence emerges as a crucial noise-reducing mechanism, allowing the system to maintain a high degree of coherence even when individuals are more prone to random changes in their state.
@@
However, escalating social feedback diminishes the extent of individual exploration, leading to a prolonged time required to reach the stationary state. This observation underscores a trade-off between speed and accuracy, indicating that a moderate level of social interactions is most beneficial when a system seeks to balance consensus accuracy and convergence time [52]. This finding aligns with previous research investigating the emergence of a speed-accuracy trade-off in opinion dynamics models [27, 41]. In those studies, interactions governed by a majority rule revealed a trade-off between decision accuracy and the size of the interaction group.
When confronted with a decision problem between two equal-quality options, the deterministic solution of the LES model predicts a system unable to break the symmetry and achieve consensus for one option. In such situations, the introduction of more complex higher-order interactions, such as cross-inhibition, becomes necessary. This need for higher-order interactions has been observed in both on-field honeybee experiments [20], computational models [23, 25, 53], and robot swarm experiments [26, 33, 54, 53]. However, finite-size fluctuations can readily break the symmetry between the options if the system is driven around the so-called ”fully interdependent” limit, where individual exploration behavior plays a minimal role. This finding suggests a simple adaptive behavior of the model: reducing the strength of the self-discovery probabilities as the decision process progresses, especially in scenarios where options are advertised similarly. By doing so, the system can break the symmetry even between identical options. However, it’s important to note that while this mechanism allows for simple symmetry breaking, it makes the system less flexible when it needs to adapt in changing environments [28, 32].
We have discovered an interesting correspondence between the behavior of the model at the fully interdependent limit and the well-known contact process [43]. In this limit, where the discovery probabilities are negligible, the dynamics are solely governed by interactions. By adjusting the interdependence, which represents the strength of these interactions, the system undergoes a non-equilibrium phase transition. Below a critical value $\tilde{\lambda}_{c}$, the system settles into an inactive, absorbing state—referred to as the uncommitted state in opinion dynamics terminology. Conversely, above the critical value, the system can maintain a non-negligible proportion of active or committed population for either of the two options, particularly when they are of equal quality. We have derived the model equations, which exactly map to the mean-field contact process equations. In addition, making use of $d=2$ lattice simulations, we have explored the critical behavior and some of the critical exponents surrounding the phase transition. Our findings confirm that the critical exponents of the order parameter and susceptibility consistently align with those reported for the contact process in $d=2$, which falls within the directed percolation universality class.
@@
We have further evaluated the model’s compliance with Weber’s Law of psychological perception, which suggests that the smallest noticeable change between two stimuli is proportional to the intensity of the base stimulus [47]. This relation was first studied in individual organisms, but recently efforts have been devoted to study swarms faced with perception tasks as superorganisms, capable of obeying the same laws [30, 36]. This assessment was conducted in both symmetric and asymmetric discovery scenarios. While the discovery probabilities do impact the discernible quality differences for a given level of interdependence, they do not alter the linear relationship with the base stimulus strength. Thus, our findings indicate a broad concordance with Weber’s Law. In this context, we have also confirmed that the system’s capability to distinguish between two similar stimuli enhances with an increase in system size.
@@
While our primary focus has been on a binary decision problem, we have also evaluated the model’s robustness when expanding the number of available options. Unlike other studies [24], we observe that increasing the number of sites consistently improves the swarm performance, or accuracy, when making decisions among non-equivalent options. As discussed in [24], the main distinction between our approaches lies in the absence or inclusion of negative social feedback, such as cross-inhibition between populations representing different options. Consequently, the effect of increasing social interactions, or signaling, has varying impacts on the system dynamics.

### v3

@@ III.2 Asymmetric discovery scenario
III.3 Weber’s Law of perception
Psychophysics explores how organisms perceive external stimuli, initially concentrating on human perception before broadening its scope to encompass other organisms across varying levels of biological complexity. There’s a proposition that a swarm engaged in decision-making tasks can be seen as a superorganism Passino et al. (2008), exhibiting cognitive traits akin to those found in vertebrate brains based on neurons. Building on recent research into honeybee-inspired decision-making models Reina et al. (2018b); Pais et al. (2013), we aim to ascertain whether the LES model also reflects such characteristics.
We particularly investigate Weber’s Law Fechner (1948), which posits that the smallest perceptible change between two stimuli (known as the just noticeable difference) is proportional to the intensity of the base stimulus. In our context, the stimuli correspond to the qualities of the two options under consideration in a binary decision problem, and the base stimulus intensity can be defined as the mean quality of the options, denoted as $\bar{q}$. According to Weber’s Law, we anticipate observing a linear correlation between the difference in quality and the base stimulus strength ($\bar{q}$), expressed as $q_{2}-q_{1}=w\bar{q}$, where $w$ represents the Weber fraction.
Following the analysis in Reina et al. (2018b), one can also relate the magnitude of finite-size fluctuations within the decision process with the random fluctuations typically observed in a discrimination process. These fluctuations can influence an organism’s ability to accurately discriminate between two stimuli Thurstone (1927); Stevens (1957). It is widely recognized that collective decisions tend to improve when made by larger groups Austen-Smith and Banks (1996); Marshall et al. (2017). As discussed in Sec. III.4, fluctuations diminish as the system size increases. Consequently, we anticipate that the system’s ability to discriminate between two similar stimuli will improve with the system size, i.e. the just noticeable difference will decrease with $N$. In agreement with Reina et al. (2018b), we observe this effect in our model, as illustrated in Fig. 7C, which also exhibits a comparable exponential trend.
In the study by Reina et al. Reina et al. (2018b), individual commitment transitions (i.e., discovery) are linked to site qualities, unlike in our model. Here, the quality-sensitivity influences system dynamics solely in the uncommitment and recruitment transitions. However, this difference does not appear to impact the model’s adherence to Weber’s Law. The independent discovery probabilities solely affect the specific value of quality differences that can be discerned. Figures 7 A and B also demonstrate that, for a given stimulus strength, the ability to discriminate decreases with $\lambda$. In other words, stronger social interaction enables the system to distinguish between more similar options, but this comes at the expense of a slower decision-making process, as previously noted in Figs. 4 and 6.
Contrary to our findings, Reina et al. in Reina et al. (2018b) observe that the noticeable quality difference increases (while the decision time decreases) with the signaling ratio, which in their setup reflects the strength of social interactions relative to individual transitions. This disparity underscores the main distinction between the two models: the presence (or absence, in our case) of negative feedback in social interactions, specifically cross-inhibition. Introducing this mechanism in the decision process leads to quicker decisions, especially in situations involving similarly or equally valued options Reina et al. (2017), albeit at the expense of reduced accuracy in the final outcome. These results emphasize not only the significance of social interactions but also their nature, as the presence or absence of cross-inhibition yields markedly different outcomes in scenarios initially featuring very similar problems.
III.4 Finite Size Effects
@@
In the lower-left corner of Fig. 8B, we observe a null variance, followed by a sudden increase at higher values of $\lambda$. This corresponds to the region where the system ends up in a fully uncommitted state, occurring when $\pi_{1,2}=0$ and $\lambda<r_{2}$ in an infinite system (recall Sec. II.1.2). Although analytically the change in behavior is predicted to occur precisely at $\lambda=r_{2}$, we note that due to finite-size effects, this behavior extends slightly beyond this threshold. At $\lambda\gtrsim r_{2}$, fluctuations suddenly increase as most of the simulation results converge around the predicted stable solution, $f_{0}=r_{2}/\lambda$, while, due to finite-size effects, some results still show the system stabilizing at the fully uncommitted state. As the system size increases, the transition becomes sharper around the threshold $\lambda=r_{2}$.
The stationary results and phase diagrams for the case where $\pi_{\alpha}\neq 0$ are independent of the initial conditions, as the system consistently evolves towards a unique stable fixed point. However, initial conditions do affect the deliberation time required to reach this stationary state. For this reason, we have chosen to analyze the system’s evolution starting from a fully uncommitted initial state, as it closely resembles a natural decision-making scenario.
At or near the fully interdependent limit ($\pi_{\alpha}=0$), the effects of initial conditions become more pronounced. As the system approaches this limit, it undergoes critical slowing down due to the phase transition at $\lambda=r_{2}$ when $\pi_{1,2}=0$ (see App A.1). In such cases, the bias introduced by initial conditions can significantly extend the decision time. Furthermore, in finite systems, the interplay between fluctuations and unfavorable initial conditions may lead the system to deviate from the stable fixed point—where the superior option is typically selected-for extended periods, particularly when $q_{1}\rightarrow q_{2}$. In the exact limit $\pi_{1,2}=0$, the system could even select the lower-quality option, with initial conditions introducing a strong bias that can significantly influence the consensus outcome.
@@
Lastly, let’s consider the scenario where the site qualities are equal (and their discovery probabilities are also equal) in the fully interdependent limit ($\pi_{\alpha}=0$). As we’ve discussed in the analytical solution of the model, detailed in Sec. II.1.2, the system undergoes a phase transition between a fully uncommitted absorbing state and an active state with finite fractions of bees advertising the available sites at $\lambda=r$. We further investigate this non-equilibrium phase transition in the following section. Here, we briefly emphasize the influence of finite-size effects under this limit. While the system may temporarily fluctuate around the stationary state predicted by the deterministic equations, finite-size effects ultimately break this tendency, and the system converges towards a consensus for only one of the options, while the other option disappears. In Figs. 9, we illustrate this phenomenon, by representing the temporal evolution of $f_{0}$, $f_{1}$ and $f_{2}$ for a binary decision problem starting from two different initial conditions. Continuous lines represent the results obtained in stochastic simulations of a finite system consisting of $N=1000$ bees, while dashed lines represent the integration of deterministic equations Eqs. (4), representing the behavior of an infinite system. We observe that fluctuations introduced by finite-size effects ultimately break the symmetry between the two available options. This can occur even when one option initially appears to be favored. The initial difference between options will also affect the likelihood of observing this switch.
@@
In real-world applications, whether in ecological systems or synthetic ones like robot swarms, these fluctuations are always expected to be relevant because most of these systems are finite. An important conclusion can be drawn from these results. In ecological systems, the ability to regulate the dissemination of information is crucial for efficient decision-making. For instance, when options have comparable qualities, it may be beneficial for the group to converge quickly towards a consensus to avoid prolonged indecision. In such cases, certain members within the group may play a pivotal role by ceasing exploration or transmitting stop exploration signals to others Seeley et al. (2012); Reina et al. (2023); March-Pons et al. (2024b). These signals would effectively reduce the parameter $\pi$ to near zero, signaling to other group members to cease exploration and focus on promoting consensus for one of the options. As we will analyze in the following section, reducing this parameter brings the system closer to the state where maximal consensus is attained. Thus, this adaptive behavior allows the group to navigate decision-making scenarios with similar options more effectively, potentially avoiding resource wastage. However, deterministic modeling approaches, may struggle to capture the dynamics of decision-making processes in real finite systems in similar circumstances.
@@
III.5 Absorbing phase transition and finite size effects in the fully interdependent limit
@@
The self-discovery probabilities $\pi$ play the role of an external field, as with non-null $\pi$ values, an inactive site can spontaneously become active. Similar to the contact process, the introduction of this external field disrupts the absorbing state and consequently the transition itself, leading the system away from criticality. However, for sufficiently small values of $\pi$, the system still obeys certain scaling laws, as discussed in Hinrichsen (2000). We have also confirmed that, for instance, the order parameter in stationary conditions is compatible with the scaling law $\rho\sim\pi^{\beta/\sigma}\sim\pi^{1/2}$, with mean-field exponent values $\beta=1$ and $\sigma=2$.
@@
Non-equilibrium phase transitions are commonly observed in biological systems Muñoz (2018). These transitions, which occur far from thermodynamic equilibrium, play a crucial role in shaping the collective behavior of living organisms. As a matter of fact, we observe that this model system achieves optimal consensus at the onset of critical behavior, when approaching the fully interdependent limit $\pi_{1,2}\rightarrow 0$ (or $\lambda\rightarrow 1$), identifying this as the optimal operating point Mora and Bialek (2011); Chialvo et al. (2020). Biological systems often incorporate adaptive mechanisms, such as quorum-dependent stop signals Seeley et al. (2012), which can modulate interdependent activity or discourage independent exploration. These mechanisms allow them to eventually operate near the critical point, once sufficient information has been gathered from the environment. The adaptive ability to operate around a critical point is particularly important for collective decisions involving equivalent options (i.e., options of the same quality and discovery probability), as symmetry can only be broken by critical fluctuations around this optimal operating point. While our current model lacks an explicit mechanism to adjust the self-discovery parameter, our parametric study reveals that operating near this critical point yields maximum consensus (see Figs. 4A,B). Developing model extensions to incorporate these features would be an interesting direction for future research.
@@ III.6 Multiple sites
Up to now, our attention has been directed towards a binary decision problem, where selecting the best option is imperative. Many collective decision making models have been concerned in generalizing their dynamics to multiple option scenarios Starnini et al. (2012); Reina et al. (2017); Vazquez et al. (2019); Ramirez et al. (2022). Particularly, previous research on this model, utilizing different parameter setups, examined a specific scenario involving five options. It’s worth noting that increasing the number of available options doesn’t qualitatively alter the model’s behavior. In a general $k$-site scenario, we still observe the same interplay between independent discovery and social feedback. At high $\lambda$ values, the system effectively identifies the best site, while at low $\lambda$ values, it settles into a multi-opinion state where each site is occupied in proportion to its quality and discovery probability. Therefore, in a broader best-of-N scenario, the specific dance frequency values and consensus transitions will rely on the $\pi_{\alpha}$ and $q_{\alpha}$ parameters defining each unique scenario. However, the overall behavioral trend of the system remains consistent across different setups.
@@
As the number of available options with the same discovery probability increases, the system is exposed to a greater amount of external information. Consequently, slightly higher values of $\lambda$ are required to achieve the consensus crossover, as evidenced by the rightward shift of the consensus curves in Fig. 12A around $Q=0$. Contrarily, when the overall level of external information ($\sum\pi_{\alpha}$) remains constant but the individual discovery probabilities $\pi_{\alpha}$ decrease as the number of options increases, as shown in Fig. 12B - the system achieves consensus more readily as the number of options increases, a result that may appear counter-intuitive compared to other studies on the best-of-N scenario with different models Reina et al. (2017). However, this outcome is a direct consequence of the decision mechanisms embedded in our model: increasing $\lambda$ strengthens the positive feedback on populations associated with the best quality, leading to the rapid elimination of options with lower qualities from consideration. Consequently, the decision process becomes nearly binary between the top two options. Decreasing the overall values of the discovery probabilities simplifies the discussion process further, facilitating consensus. Supplementary Fig. 1 in sup illustrates all dance frequencies $f_{\alpha}$ in the reported scenarios, demonstrating the consistently decreasing trends of the remaining populations in favor of the highest-quality option as $\lambda$ increases.
Other scenarios, such as a generally asymmetric discovery scenario (where $\pi_{1}>\pi_{2}>...>\pi_{k}$) or equivalent inferior options ($q_{1}=...=q_{k-1}$), exhibit similar trends to the discussed symmetric discovery scenario. Naturally, the specific values of $f_{k}$ or the consensus may vary, as it becomes more challenging for the swarm to achieve consensus under these conditions. However, the interplay between interdependence, site qualities, and discovery probabilities remains unchanged. In Supplementary Fig. 2 in sup we present the results obtained in these scenarios to emphasize the qualitative similarities discussed earlier.
@@ IV Conclusions
We have performed an analytical study of an agent-based model inspired by the collective behavior of honeybees List et al. (2008), following the mathematical framework outlined in Galla (2010). This model incorporates key features such as the balance between independent exploration and social interdependence, as well as the sensitivity of agents to option quality. Using analytical reductions of this agent-based model, we extensively analyzed its behavior in the space of relevant parameters. Our goal was to investigate whether the system achieves consensus for the best available option when balancing information acquisition and social interactions. Furthermore, we supplemented this analysis with simulations to report the time required for the system to converge to its stationary state.
@@
We have explored two scenarios: one where options are equally likely to be discovered, and another where the worst quality option (among two) is more likely to be discovered. In both cases, when the system prioritizes interdependence, the established consensus is stronger. This highlights the role of social interactions in achieving a higher accuracy in the final decision, consistent with similar honeybee-inspired models Reina et al. (2015, 2017); Gray et al. (2018). Interdependence plays a critical role, especially when the system deals with high values of the self-discovery probabilities or when the superior option faces a disadvantage in terms of discovery likelihood. In such scenarios, interdependence emerges as a crucial noise-reducing mechanism, allowing the system to maintain a high degree of coherence even when individuals are more prone to random changes in their state.
@@
However, escalating social feedback diminishes the extent of individual exploration, leading to a prolonged time required to reach the stationary state. This observation underscores a trade-off between speed and accuracy, indicating that a moderate level of social interactions is most beneficial when a system seeks to balance consensus accuracy and convergence time Passino and Seeley (2006). This finding aligns with previous research investigating the emergence of a speed-accuracy trade-off in opinion dynamics models Valentini et al. (2016); Reina et al. (2024). In those studies, interactions governed by a majority rule revealed a trade-off between decision accuracy and the size of the interaction group.
@@
In our binary decision task analysis, we have further proved the model’s compliance with Weber’s Law of psychological perception, which suggests that the smallest noticeable change between two stimuli is proportional to the intensity of the base stimulus Fechner (1948). This relation was first studied in individual organisms, but recently efforts have been devoted to study swarms faced with perception tasks as superorganisms, capable of obeying the same laws Passino et al. (2008); Reina et al. (2018b). This assessment was conducted in both symmetric and asymmetric discovery scenarios. While the discovery probabilities do impact the discernible quality differences for a given level of interdependence, they do not alter the linear relationship with the base stimulus strength. Thus, our findings indicate a broad concordance with Weber’s Law. In this context, we have also confirmed that the system’s capability to distinguish between two similar stimuli enhances with an increase in system size.
When confronted with a decision problem between two equal-quality options, the deterministic solution of the LES model predicts a system unable to break the symmetry and achieve consensus for one option. In such situations, the introduction of more complex higher-order interactions, such as cross-inhibition, becomes necessary. This need for higher-order interactions has been observed in both on-field honeybee experiments Seeley et al. (2012), computational models Reina et al. (2015); Gray et al. (2018); Reina et al. (2023), and robot swarm experiments Reina et al. (2018a); Zakir et al. (2022); Talamali et al. (2019); Reina et al. (2023). However, finite-size fluctuations can readily break the symmetry between the options if the system is driven around the so-called ”fully interdependent” limit, where individual exploration behavior plays a minimal role. This finding suggests a simple adaptive behavior of the model: reducing the strength of the self-discovery probabilities as the decision process progresses, especially in scenarios where options are advertised similarly. By doing so, the system can break the symmetry even between identical options. However, it’s important to note that while this mechanism allows for simple symmetry breaking, it makes the system less flexible when it needs to adapt in changing environments Talamali et al. (2021); Aust et al. (2022).
Delving into the analysis of the fully interdependent limit, we have discovered an interesting correspondence with the well-known contact process Hinrichsen (2000). In this limit, where the discovery probabilities are negligible, the dynamics are solely governed by interactions. By adjusting the interdependence, which represents the strength of these interactions, the system undergoes a non-equilibrium phase transition. Below a critical value $\tilde{\lambda}_{c}$, the system settles into an inactive, absorbing state—referred to as the uncommitted state in opinion dynamics terminology. Conversely, above the critical value, the system can maintain a non-negligible proportion of active or committed population for either of the two options, particularly when they are of equal quality. We have derived the model equations, which exactly map to the mean-field contact process equations. In addition, making use of $d=2$ lattice simulations, we have explored the critical behavior and some of the critical exponents surrounding the phase transition. Our findings confirm that the critical exponents of the order parameter and susceptibility consistently align with those reported for the contact process in $d=2$, which falls within the directed percolation universality class. Our study further reveals that the system achieves maximum consensus near this critical point. This finding aligns with the ongoing debate on how biological and bio-inspired systems operate near criticality Mora and Bialek (2011); Muñoz (2018); Chialvo et al. (2020). It reinforces the idea that adaptive mechanisms, such as inhibitory signals, may play a crucial role in reducing independent information gathering, allowing the system to rely more on intercommunication to reach maximum consensus. These mechanisms can guide the system to operate near this critical point, once enough environmental information has been pooled, optimizing decision-making efficiency.
@@
While our primary focus has been on a binary decision problem, we have also evaluated the model’s robustness when expanding the number of available options. Unlike other studies Reina et al. (2017), we observe that increasing the number of sites consistently improves the swarm performance, or accuracy, when making decisions among non-equivalent options. As discussed in Reina et al. (2017), the main distinction between our approaches lies in the absence or inclusion of negative social feedback, such as cross-inhibition between populations representing different options. Consequently, the effect of increasing social interactions, or signaling, has varying impacts on the system dynamics.